OpenStudy (anonymous):

differential equation (D3-11D-20)y=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

is it (d^3-11d-20)y=0 ?

OpenStudy (anonymous):

differential equation (D3-11D-20)y=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

differential equation (D-11D-20)y=0

OpenStudy (anonymous):

what is D here?

OpenStudy (anonymous):

write it please with dy/dx notation

OpenStudy (anonymous):

upper of D is 3

OpenStudy (anonymous):

differential equation (D-11D-20)y=0 please help

OpenStudy (anonymous):

I think I dont know it than. sorry

OpenStudy (anonymous):

differential equation (D-11D-20)y=0

OpenStudy (anonymous):

Please Help Me